OpenStudy (anonymous):

If the gravitational constant is 20, and the initial velocity is 40, what is the maximum height?

OpenStudy (anonymous):

use the kinematic equation- V^2 = V^2 + 2ah so 0^2 = 40^2 +2(-20)*h make sense?